OpenStudy (jazzyfa30):

Find the exact value of the expension. sin(7pi/24)cos (pi/8)-cos(7pi/24)sin(pi/8) A)(squareroot2)/2 B)1 C)1/2 D)(squareroot3)/2 @ganeshie8

ganeshie8 (ganeshie8):

hint : sin(A-B)

ganeshie8 (ganeshie8):

\[\sin (A-B) = \sin A \cos B - \cos A \sin B\]

ganeshie8 (ganeshie8):

you need to apply it in reverse..

OpenStudy (jazzyfa30):

???????

ganeshie8 (ganeshie8):

\[\sin(7\pi/24)\cos (\pi/8)-\cos(7\pi/24)\sin(\pi/8) = \sin (7\pi/24 - \pi/8 ) \]

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

\[\sin(7\pi/24)\cos (\pi/8)-\cos(7\pi/24)\sin(\pi/8) = \sin (7\pi/24 - \pi/8 ) \\ = \sin (4 \pi/24) = \sin (\pi / 6) = ?\]

OpenStudy (jazzyfa30):

1/2

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark }\)

OpenStudy (jazzyfa30):

is that the answer

OpenStudy (jazzyfa30):

ok 2 more and we are done thank you so much

ganeshie8 (ganeshie8):

np :)